Find the Product Help Please!

-xy^5(2xy^2+7xy-4x)

Answer by checkley75(3666)

-xy^5(2xy^2+7xy-4x)
-xy^5*2xy^2=-2x^2y^7
-xy^5*7xy=-7x^2y^6
-xy^5*-4x=4x^2y^5 thus we have the following
-2x^2y^7-7x^2y^6+4x^2y^5